OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (anonymous):

The slope is the same so use the equation y = x + b y = x + b Now plug in the coordinate and solve for b. 10 = 4 + b b = ? y = x + ?